# Turán number of complete multipartite graphs in multipartite graphs

### v1

@@ 3. $(R3)$
$\sum_{i=2}^{r}e_{G}(\tilde{U}_{i})$ $\leq\sum_{i\in L_{1}}(t-1)\min\{z_{i},w_{i}\}+\sum_{i\in L_{2}}((t-1)w_{i}+\alpha_{i})+\sum_{i\in L_{3}}((t-1)(z_{i}-n+w_{i})+\alpha_{i})$
$=(t-1)(bn-w_{1}+n-z_{1})+\sum_{i\in L_{1}}(t-1)\min\{n-w_{i},n-z_{i}\}+\sum_{i\in L_{2}\cup L_{3}}\alpha_{i}.$
Recall that $e(G^{\prime})\leq t_{r}(k)n^{2}+\sum_{i\geq 2}(\beta_{i}-\alpha_{i})$ by Claim 4.4 and (4.16). For $i\in L_{1}\cup L_{2}$, we have $\tilde{Z}_{i}=Z^{i}_{i}$ and thus $\beta_{i}=0$ by (4.25). It follows that (note that $L\cap L_{3}=\emptyset$)
$\sum_{i\geq 2}\beta_{i}=\sum_{i\in L_{3}}\beta_{i}=\sum_{i\in L_{3}}\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|W_{j}^{\prime}|-n)=\sum_{j\in L}\sum_{i\in L_{3}\setminus\{j\}}|Z_{j}^{i}|(|W_{j}^{\prime}|-n)=\sum_{j\in L}(n-z_{j})(w_{j}-n)$
Note that $1\in L$ by (4.19) and $(n-z_{1})(w_{1}-n)\leq 0$ by Claim 4.8. Furthermore, since $n-z_{j}=0$ for $j\in L_{2}$, it follows that $\sum_{i\geq 2}\beta_{i}=\sum_{j\in L_{1}}(n-z_{j})(w_{j}-n)$. Consequently, $e(G^{\prime})\leq t_{r}(k)n^{2}+\sum_{j\in L_{1}}(n-z_{j})(w_{j}-n)-\sum_{i\geq 2}\alpha_{i}$.
Recall that $e_{G}(\tilde{U}_{1})=f_{1}\leq z_{t}(z_{1},w_{1})$. By ($E3$), we have $z_{t}(z_{1},w_{1})+(t-1)(n-z_{1}+n-w_{1})\leq z_{t}(n)$. Thus, combining these estimates together, we get
$e(G)=e(G^{\prime})+\sum_{i=1}^{r}e_{G}(\tilde{U}_{i})\leq t_{r}(k)n^{2}+z_{t}(n)+(t-1)(b-1)n+y,$
@@
$(t-1)\min\{n-w_{i},n-z_{i}\}-(n-z_{i})(n-w_{i})=y_{i}(t-1-y_{i}^{\prime})\leq\lfloor(t-1)^{2}/4\rfloor.$
Since $L_{1}\subseteq L\setminus\{1\}$, we have $|L_{1}|\leq b-1$. Consequently, $e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+(t-1)(b-1)n+(b-1)\lfloor(t-1)^{2}/4\rfloor=g_{2}(n,r,k,t)$. ∎

### v2

@@ 3. $(R3)$
$\sum_{i=2}^{r}(e_{G}(\tilde{U}_{i})-\alpha_{i})$ $\leq\sum_{i\in L_{1}}(t-1)\min\{z_{i},w_{i}\}+\sum_{i\in L_{2}\cup L_{3}}(t-1)(z_{i}-n+w_{i})-\lambda$
$=(t-1)(bn-w_{1}+n-z_{1})+\sum_{i\in L_{1}}(t-1)\min\{n-w_{i},n-z_{i}\}-\lambda.$
Finally, we work on the $\beta_{i}$’s and recall that $\beta_{i}=\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}|-n+|Z_{i}\setminus\tilde{Z}_{i}|)$. For $i\in\{1\}\cup L_{1}\cup L_{2}$, as $\tilde{Z}_{i}\setminus Z_{i}=\emptyset$, $\beta_{i}=0$. For $i\in L_{3}$, as $|W_{i}|<2\varepsilon n$, we have $Z_{i}\setminus\tilde{Z}_{i}=\emptyset$; for any $j\in L\setminus\{i\}$, we have $\tilde{Z}_{j}\setminus Z_{j}=\emptyset$ and $W^{\prime}_{j}\subseteq V_{q_{j}}$ again by Claim 4.10. Hence $\beta_{i}=\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|W_{j}^{\prime}|-n)\leq 0$ because $|W^{\prime}_{j}|\leq n$. It then follows that (noting that $L\cap L_{3}=\emptyset$)
$\sum_{i\geq 1}\beta_{i}$ $=\sum_{i\in L_{3}}\beta_{i}=\sum_{i\in L_{3}}\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|W_{j}^{\prime}|-n)$
$=\sum_{j\in L}\sum_{i\in L_{3}\setminus\{j\}}|Z_{j}^{i}|(|W_{j}^{\prime}|-n)=\sum_{j\in L}(n-z_{j})(w_{j}-n).$
Note that $1\in L$ by (4.16) and $(n-z_{1})(w_{1}-n)\leq 0$ by Claim 4.8. Furthermore, since $n-z_{j}=0$ for $j\in L_{2}$, it follows that $\sum_{i\geq 1}\beta_{i}=\sum_{j\in L_{1}}(n-z_{j})(w_{j}-n)$.
Recall that $e_{G}(\tilde{U}_{1})=f_{1}\leq z_{t}(z_{1},w_{1})$. By ($E3$), we have $z_{t}(z_{1},w_{1})+(t-1)(n-z_{1}+n-w_{1})\leq z_{t}(n)$. Thus, combining these estimates together, by (4.7), we get
$e(G)\leq t_{r}(k)n^{2}+\sum_{i\in[r]}(e_{G}(\tilde{U}_{i})-\alpha_{i}+\beta_{i})\leq t_{r}(k)n^{2}+z_{t}(n)+(t-1)(b-1)n+y-\lambda,$ (4.21)
@@
$(t-1)\min\{n-w_{i},n-z_{i}\}-(n-z_{i})(n-w_{i})=y_{i}(t-1-y_{i}^{\prime})\leq\lfloor(t-1)^{2}/4\rfloor\leq 1.$
Since $L_{1}\subseteq L\setminus\{1\}$, we have $|L_{1}|\leq b-1$. Moreover, by Claim 4.10, we have $w_{i}\leq|W_{i}|+|W_{i}^{\prime}\setminus W_{i}|\leq 2\varepsilon n+\gamma n\leq 3\varepsilon n$ for each $i\in L_{3}$. If $|L_{1}\cup L_{2}|\leq b-2$, then
$bn=\sum_{i\in[r]}|W_{i}|\leq n+(b-2)n+(r-b+1)\cdot 3\varepsilon n<bn,$
a contradiction. This implies $|L_{1}\cup L_{2}|\geq b-1$, and $|L_{3}|\leq r-b$. Since $|L_{1}|-\lambda=\min\{|L_{1}|,|L_{3}|\}\leq|L_{3}|\leq r-b$, it follows that $|L_{1}|-\lambda\leq\min\{b-1,r-b\}$. Consequently, as $\lfloor(t-1)^{2}/4\rfloor\leq 1$, we get
$y-\lambda\leq|L_{1}|\lfloor(t-1)^{2}/4\rfloor-\lambda\leq\min\{b-1,r-b\}\lfloor(t-1)^{2}/4\rfloor,$
Together with (4.21), it gives the desired bound $e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+(t-1)(b-1)n+\min\{b-1,r-b\}\lfloor(t-1)^{2}/4\rfloor=g(n,r,k,t)$. This completes the proof of Theorem 3 for $k<2r$. ∎
Acknowledgments
The authors would like to thank Junxue Zhang for creating a figure and helpful comments.